OpenStudy (anonymous):

Generate the first 5 terms of this sequence: f(1) = 0 and f(2) = 1, f(n) = f(n - 1) + f(n - 2), for n > 2. 0, -1, 1, 0, 2 0, 1, 1, 2, 3 0, 1, 2, 2, 3 0, 1, 1, 2, 2

terenzreignz (terenzreignz):

This is a rather well known sequence :D \[\Large \left.\begin{matrix}f(1)&f(2)&f(3)&f(4)&f(5)\\0&1&?&?&?\end{matrix}\right.\]

terenzreignz (terenzreignz):

f(n) = f(n-1) + f(n-2) So what happens when n = 3?

OpenStudy (anonymous):

n=2

OpenStudy (anonymous):

it would be 2 3 4 after that ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

would the answer be C

terenzreignz (terenzreignz):

I don't know, what's f(3)? Use the fact that f(n) = f(n-1) + f(n-2)

OpenStudy (anonymous):

n-1 would be 2 and n-2 would be 1?

terenzreignz (terenzreignz):

That's good :) so f(3) = f(2) + f(1) Which is...?

OpenStudy (anonymous):

6?

terenzreignz (terenzreignz):

Don't be silly :P You're already given the values of f(2) and f(1)

terenzreignz (terenzreignz):

\[\Large \left.\begin{matrix}\color{blue}{f(1)}&\color{blue}{f(2)}&f(3)&f(4)&f(5)\\\color{red}0&\color{red}1&?&?&?\end{matrix}\right.\]

OpenStudy (anonymous):

0 1 2 2 3?

terenzreignz (terenzreignz):

I don't know, let's find out together, but step-by-step. Now, what's f(3) = ?

OpenStudy (anonymous):

2

terenzreignz (terenzreignz):

How did you get 2?

OpenStudy (anonymous):

just going 1 number less than 3

OpenStudy (anonymous):

since it looks like the others did that

terenzreignz (terenzreignz):

I told you to follow this formula (as it is said in your problem) f(n) = f(n-1) + f(n-2) Set n = 3, and find f(3)

OpenStudy (anonymous):

i got 3

terenzreignz (terenzreignz):

How?

OpenStudy (anonymous):

plugging in 3 where n is

terenzreignz (terenzreignz):

Okay... so how did you get 3?

terenzreignz (terenzreignz):

I don't see how you could get 3 by plugging in n = 3...

OpenStudy (anonymous):

(3-1)=2 + (3-2)=1 and i just added it to 3 i dont get it

terenzreignz (terenzreignz):

f(n) = f(n-1) + f(n-2) Please plug in n = 3 again, but be very careful when evaluating.

OpenStudy (anonymous):

what did i do wrong

terenzreignz (terenzreignz):

It's not (3-1) and (3-2) which you add, it's f(3-1) and f(3-2) which are totally different.

OpenStudy (anonymous):

so what do i do with the f

OpenStudy (anonymous):

you multiply?

terenzreignz (terenzreignz):

It's a function. And as I have said before, you ARE given the values of f(1) and f(2)

OpenStudy (anonymous):

i think its C

terenzreignz (terenzreignz):

But why?

OpenStudy (anonymous):

because it looks like thats how the pattern goes am i wrong

OpenStudy (anonymous):

follow pony

terenzreignz (terenzreignz):

I strongly advise against guessing ^_^

OpenStudy (anonymous):

is my guess wrong

terenzreignz (terenzreignz):

Let's find out. Starting with you figuring out f(3) from the formula f(n) = f(n-1) + f(n-2)

OpenStudy (anonymous):

f(2)=1

OpenStudy (anonymous):

f(1)=0

OpenStudy (anonymous):

f(3)=2

OpenStudy (anonymous):

so 1+0=1

OpenStudy (anonymous):

what do you think

OpenStudy (anonymous):

what do i think is the answer?

terenzreignz (terenzreignz):

Just f(3) for now. How do you get 2?

OpenStudy (anonymous):

3-1=2

terenzreignz (terenzreignz):

But why would you do 3-1?

OpenStudy (anonymous):

n-1 u plug in the 3

terenzreignz (terenzreignz):

but n-1 isn't the expression to plug n = 3 in It's f(n) = f(n-1) + f(n-2)